# Bank–Laine functions with preassigned number of zeros

### v2

@@ 1. Introduction
Let $n\in\mathbb{N}$ be a positive integer and $\lambda_{1},\lambda_{2}\in[0,n]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. Then there exists an entire function $A(z)$ of order $n$ such that the Bank–Laine function $E=f_{1}f_{2}$ for two linearly independent solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda_{1}$ and $\log n(r,0,f_{2})\asymp(\log r)^{2}$ when $\lambda_{2}=0$ or $\log n(r,0,f_{2})\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq n$. Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
A question concerning the coefficient function $A(z)$ of the second order differential equation (1.1) was asked in [8, Question 4.5]: Is it possible to characterize the transcendental entire functions $A(z)$ where (1.1) admits a non-trivial solution $f_{1}$ with $\lambda(f_{1})<\infty$, such that for every solution $f_{2}$ of (1.1) that is linearly independent with $f_{1}$ we have $\lambda(f_{2})=\infty$? The construction in Theorem 1.6 and Theorem 1.7 yields such entire functions $A(z)$ of finite order. With modifications of the methods there, we may construct more such coefficient functions $A(z)$ with certain growth properties.
In the remainder of this paper, we shall provide the proof for Theorem 1.3, Theorem 1.4, Theorem 1.6 and Theorem 1.7 in Section 2, Section 3, Section 4 and Section 5, respectively. In the following sections, we keep the permanent notation: $(f_{1},f_{2})$ is a normalized pair of solutions of (1.1), the quotient $F=f_{2}/f_{1}$ is a solution of $S(F)=2A$, and $E=f_{1}f_{2}=F/F^{\prime}$ is a solution of (1.8).
@@
For every pair $(m,2n)$ of two nonnegative integers $m$ and $n$, we consider the rational function $R_{m,n}(z)$ in (1.4). In this section, we fix two distinct pairs $(m_{1},2n_{1})$ and $(m_{2},2n_{2})$ and will sometimes omit them from notation. We shall assume that at least one of $m_{1}$ and $m_{2}$ is nonzero and at least one of $n_{1}$ and $n_{2}$ is nonzero. Note that, unlike in [4], $g_{m_{1},n_{1}}$ and $g_{m_{2},n_{2}}$ are locally univalent meromorphic functions. The process of gluing two different $g_{m,n}$s is almost the same as that in [4]: by restricting $g_{m_{1},n_{1}}$ to the lower half-plane $H^{-}$ and $g_{m_{2},n_{2}}$ to the upper half-plane $H^{+}$, we then paste them together, using a quasiconformal surgery, producing a meromorphic function $F$. Then our Bank–Laine function will be $E=F/F^{\prime}$ and thus $A=B(E)/4$ as in (1.8).
@@
The asymptotic behaviors of the diffeomorphism $\phi$ satisft:
@@
$\begin{split}c=\frac{1}{M}\log\left[\frac{\binom{m_{1}+2n_{1}}{m_{1}}}{\binom{m_{2}+2n_{2}}{m_{2}}}\frac{M!}{N!}\right]\quad\text{and}\quad\delta=\frac{1}{2}\min\{1,K\}.\end{split}$
@@
$\begin{split}\log g_{m_{1},n_{1}}(\phi(x))&=e^{\phi(x)}(1+O(\phi(x)e^{-\phi(x)}))\\ &=e^{\phi(x)}(1+O(\phi(x)e^{-2x/3})),\quad x\to\infty.\end{split}$
@@
(2.4) $\begin{split}\phi^{\prime}=\frac{g^{\prime}_{m_{2},n_{2}}}{g_{m_{2},n_{2}}}\frac{g_{m_{1},n_{1}}\circ\phi}{g^{\prime}_{m_{1},n_{1}}\circ\phi}=e^{Nx-M\phi(x)}\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}\frac{\sum_{i=0}^{m_{1}}A_{i}e^{i\phi(x)}}{\sum_{i=0}^{m_{2}}A_{i}e^{ix}}\frac{\sum_{j=0}^{2n_{1}}B_{j}e^{j\phi(x)}}{\sum_{j=0}^{2n_{2}}B_{j}e^{jx}},\end{split}$
@@
(2.5) $\begin{split}\phi^{\prime}&=e^{Nx-M\phi(x)+(m_{1}+2n_{1})\phi(x)-(m_{2}+2n_{2})x}\frac{1+O(e^{-\phi(x)})}{1+O(e^{-x})}\frac{1+O(e^{-\phi(x)})}{1+O(e^{-x})}\\ &=e^{O(e^{-x/2})}(1+O(e^{-\phi(x)})+O(e^{-x}))=1+O(e^{-x/2}),\quad x\to\infty.\end{split}$
@@
$\begin{split}\frac{\binom{m_{2}+2n_{2}}{m_{2}}}{\binom{m_{1}+2n_{1}}{m_{1}}}\frac{N!}{M!}e^{M\phi(x)-Nx}=1+O(e^{x})+O(e^{\phi(x)}),\quad x\to-\infty\end{split}$
@@
$\begin{split}\phi(x)=\frac{N}{M}x+\frac{1}{M}\log\left[\frac{\binom{m_{1}+2n_{1}}{m_{1}}}{\binom{m_{2}+2n_{2}}{m_{2}}}\frac{M!}{N!}\right]+O(e^{x})+O(e^{\phi(x)}),\quad x\to-\infty,\end{split}$
@@
$\begin{split}\phi^{\prime}(x)&=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{Nx-M\phi(x)}\frac{1+O(e^{\phi(x)})}{1+O(e^{x})}\frac{1+O(e^{\phi(x)})}{1+O(e^{x})}\\ &=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{Nx-M(Kx+c+O(e^{-\delta|x|/2}))}(1+O(e^{\phi(x)})+O(e^{x}))\\ &=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{-Nc+O(e^{-\delta|x|/2})}(1+O(e^{\phi(x)})+O(e^{x}))=K+O(e^{-\delta|x|/2}).\end{split}$
Now we begin to prove Theorem 1.3. We shall first define a quasiregular function in the same way as in [5]. Though the functions $g_{m_{1},n_{1}}$ and $g_{m_{2},n_{2}}$ we use here may have some poles, the process of definition here is in many parts almost the same as that in [5]. For the convenience of calculating the number of the zeros of the desired Bank–Laine functions in the final part of the proof and also of the applications in the proof of Theorem 1.6 in Section 4 and the proof of Theorem 1.7 in Section 5, below we shall keep as many as details.
@@
We note that it is possible that $K=1$ with our choice of $(m_{1},n_{1})$ and $(m_{2},n_{2})$. As shown in [5], $p$ maps $D$ onto the complement $G$ of a logarithmic spiral $\Gamma$, with
@@
In particular, when $K=1$, $p$ is just the identity in the plane and $\Gamma=\mathbb{R}_{\leq 0}$. The inverse map $h=p^{-1}$ is a conformal homeomorphism $h:G\to D$. Let $\Gamma^{\prime}=p(\mathbb{R}_{\geq 0})$. The two logarithmic spirals $\Gamma$ and $\Gamma^{\prime}$ divide the plane into two parts $G^{+}$ and $G^{-}$, which are images under $p$ of the upper and lower half-planes, respectively.
@@
$V(z)=\left\{\begin{array}[]{ll}(g_{m_{2},n_{2}}\circ h)(z),&\quad z\in G^{+},\\ (g_{m_{1},n_{1}}\circ h)(z),&\quad z\in G^{-}\end{array}\right.$
is meromorphic in $G^{+}\cup G^{-}$ and has a jump discontinuity on $\Gamma$ and $\Gamma^{\prime}$. To remove this discontinuity, we consider the strip $\Pi=\{z:|\text{Im}\ z|<1\}$ and define a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ to be the identity outside of $\Pi$ and for $y=\text{Im}\ z\in(-1,1)$,
(2.8) $\tau(x+iy)=\left\{\begin{array}[]{ll}\phi(x)+|y|(x-\phi(x))+iy,&\quad x\geq 0,\\ \phi(x/K)+|y|(x-\phi(x/K)+iy),&\quad x<0.\end{array}\right.$
Then $\tau$ commutes the complex conjugation, and satisfies
(2.9) $\begin{split}\tau(x)=\phi(x),\quad x>0,\quad\text{and}\quad\tau(Kx)=\phi(x),\quad x<0.\end{split}$
The Jacobian matrix $D_{\tau}$ of $\tau$ is given for $x>0$ and $10<|y|<$ 1 by
$D_{\tau}(x+iy)=\left(\begin{array}[]{cc}\phi^{\prime}(x)+|y|(1-\phi(x))&\pm(x-\phi(x))\\ 0&1\\ \end{array}\right),$
@@
$D_{\tau}(x+iy)=\left(\begin{array}[]{cc}1&0\\ 0&1\\ \end{array}\right),\quad 0<|y|<1,\quad x\to\infty.$
@@
$D_{\tau}(x+iy)=\left(\begin{array}[]{cc}1&\mp c\\ 0&1\\ \end{array}\right),\quad 0<|y|<1,\quad x\to-\infty.$
We conclude that $\tau$ is quasiconformal in the plane.
@@
(2.10) $U(z)=\left\{\begin{array}[]{ll}(g_{m_{2},n_{2}}\circ h)(z),&\quad z\in G^{+}\cup\Gamma\cup\Gamma^{\prime}\cup\{0\},\\ (g_{m_{1},n_{1}}\circ\tau\circ h)(z),&\quad z\in G^{-}.\end{array}\right.$
@@
It follows from (2.7) and (2.8) that $U$ is continuous on $\Gamma$ and $\Gamma^{\prime}$ and quasiregular in the plane. The existence theorem for solutions of the Beltrami equation [13, § V.1] shows that there exists a quasiconformal homeomorphism $\psi:\mathbb{C}\to\mathbb{C}$ with the same Beltrami coefficient as $U$. The function $F=U\circ\psi^{-1}$ is then meromorphic. Moreover, as shown in [4], the Beltrami coefficient of $U$ (and hence of $\psi$) satisfies the hypotheses of the Techmüler–Wittich–Beliskii theorem [13, § V.6] (see also [16]). This theorem shows that $\psi$ is conformal at $\infty$ and may thus be normalized to satisfy
(2.11) $\begin{split}\psi(z)\sim z,\quad z\to\infty.\end{split}$
@@
We set $E=F/F^{\prime}$. As $F^{\prime}(z)\not=0$ for all $z\in\mathbb{C}$ by construction, $E$ is entire. As all zeros and poles of $F$ are simple, all residues of $F/F^{\prime}$ are equal to $1$ or $-1$, so $E^{\prime}(z)=1$ or $E^{\prime}(z)=-1$ at every zero of $E$, which implies the Bank–Laine property. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A(z)$.
@@
We shall use a different method from that in [4] to show that $E$ and also the entire function $A(z)$ both have order $\rho=1/\text{Re}(\mu)$. Below we shall first estimate the numbers of poles and zeros of $F$ respectively. For two positive quantities $S_{1}(r)$ and $S_{2}(r)$, below we use the notation $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$.
@@
For the meromorphic function $g_{m_{2},n_{2}}$, if $m_{2}\not=0$, then the number of poles of $g_{m_{2},n_{2}}$ satisfies $n(r,\infty,g_{m_{2},n_{2}})\sim m_{2}r/\pi$ as $r\to\infty$. Thus, by the periodicity of $g_{m_{2},n_{2}}$, the number of poles of $g_{m_{2},n_{2}}$ in the upper half-pane $H^{+}$ satisfies $\hat{n}(r,\infty,g_{m_{2},n_{2}})\sim m_{2}r/(2\pi)$ as $r\to\infty$. We note that the image of the circle $\{z:|z|=r\}$ under $h(z)=z^{1/\mu}$ is the part of the logarithmic spiral which connects two points on the negative real axis and intersects the positive real axis at $r^{1/\text{Re}(\mu)}$. By the definitions of $p$ and $h$ we may write $z=re^{i\theta}$ on the circle such that $z_{0}=re^{i\theta_{0}}$ is a point on $\Gamma^{\prime}$ and
$\begin{split}h(z)=\exp\left(\frac{1}{\mu}\log z\right)=\exp\left(\frac{[\text{Im}(\mu)]\theta+[\text{Re}(\mu)]\log r}{|\mu|^{2}}+i\frac{[\text{Re}(\mu)]\theta-[\text{Im}(\mu)]\log r}{|\mu|^{2}}\right).\end{split}$
@@
When $K>1$, noting that $\text{Re}(\mu)>0$ and $\text{Im}(\mu)<0$, we may write $-(\theta_{0}+\pi)\leq\theta\leq-(\theta_{0}-\pi)$, where $\theta_{0}$ satisfies $(\text{Re}(\mu))\theta_{0}-(\text{Im}(\mu))\log r=0$. A simple computation shows that
@@
Therefore, when $r$ is large, the number of poles of $g_{m_{2},n_{2}}\circ h$ in $G^{+}$ satisfies
$\begin{split}c_{2}r^{1/\text{Re}(\mu)}\leq\hat{n}(r,\infty,g_{m_{2},n_{2}}\circ h)\leq c_{1}r^{1/\text{Re}(\mu)}\end{split}$
for two positive constants $c_{1}$ and $c_{2}$ both dependent on $K$. Similarly, if $m_{1}\not=0$, the number of poles of $g_{m_{1},n_{1}}\circ h$ in $G^{-}$ satisfies
(2.13) $\begin{split}c_{4}r^{1/\text{Re}(\mu)}\leq\check{n}(r,\infty,g_{m_{1},n_{1}}\circ h)\leq c_{3}r^{1/\text{Re}(\mu)}\end{split}$
for two positive constants $c_{3}$ and $c_{4}$ both dependent on $K$. By the definition of $\tau(x+iy)$ in (2.8) and the asymptotic behaviors for $\phi$ in (2.2) and (2.3), we see that the number of poles of $g_{m_{1},n_{1}}\circ\tau\circ h$ in $G^{-}$ satisfies
(2.14) $\begin{split}c_{6}r^{1/\text{Re}(\mu)}\leq\check{n}(r,\infty,g_{m_{1},n_{1}}\circ\tau\circ h)\leq c_{5}r^{1/\text{Re}(\mu)}\end{split}$
for two positive constants $c_{5}$ and $c_{6}$ both dependent on $K$. When $K<1$ or $K=1$, we can also obtain two similar estimates as in (2.13) and (2.14). By summarizing the above estimates together with the relation in (2.11), we conclude that $n(r,\infty,F)\asymp r^{1/\text{Re}(\mu)}$. It follows that $N(r,\infty,F)\asymp r^{1/\text{Re}(\mu)}$. Similarly, when at least one of $n_{1}$ and $n_{2}$ is nonzero, for the zeros of $F$ we also obtain $n(r,0,F)\asymp r^{1/\text{Re}(\mu)}$ and thus $N(r,0,F)\asymp r^{1/\text{Re}(\mu)}$. Since $E=F/F^{\prime}$ and since the poles and zeros of $F$ are simple, we have
(2.15) $\begin{split}N\left(r,0,E\right)\asymp r^{1/\text{Re}(\mu)}.\end{split}$
@@
where $c$ is a suitably chosen positive constant and, together with the asymptotic behaviors for $\phi$ in (2.2) and (2.3) and also the definition $\tau(x+iy)$ in (2.9), that
$\begin{split}\log|g_{m_{1},n_{1}}\circ\tau\circ h|\leq c(L_{1}+2\varepsilon)|z|^{1/\text{Re}(\mu)}+\exp(c|z|^{1/\text{Re}(\mu)})\quad\text{for}\quad z\in G^{-}\quad\text{and}\quad|z|\not\in\Omega_{3}.\end{split}$
@@
as $|z|\to\infty$ outside a set $\Omega_{3}$ of finite linear measure. Together with the relation in (2.17), the lemma on the logarithmic derivative now implies that $E=F/F^{\prime}$ satisfies
@@
outside a set of finite linear measure. Thus, by (2.15) and the first main theorem of Nevanlinna theory [9, Chapter 1], the characteristic function $T(r,E)$ of $E$ satisfies
$\begin{split}T\left(r,E\right)\asymp r^{1/\text{Re}(\mu)}.\end{split}$

### v3

@@ 1. Introduction
Let $n\in\mathbb{N}$ be a positive integer and $\lambda_{1},\lambda_{2}\in[0,n]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$ and $\lambda_{3}\in[0,\infty)$. Then there exists an entire function $A$ of order $n$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda_{1}$ and $n(r,0,f_{2})\asymp r^{\lambda_{3}}$ or $\log n(r,0,f_{2})\asymp(\log r)^{2}$ when $\lambda_{2}=0$ or $\log n(r,0,f_{2})\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq n$. Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
We note that the Bank–Laine functions constructed in Theorem 1.6 and Theorem 1.7 all have infinite order. However, in Theorem 1.7 where $\rho(A)$ is an integer, each of the three cases $\lambda(E)>\rho(A)$, $\lambda(E)=\rho(A)$ or $\lambda(E)<\rho(A)$ can occur. Together with Theorem 1.4, this shows that the Bank–Laine function $E$ associated with an entire coefficient $A$ of integer order indeed have value distribution properties not covered by Theorem 1.1.
A question concerning the entire function $A$ in the second order differential equation (1.1) was asked in [9, Question 4.5]: Is it possible to characterize the transcendental entire functions $A$ where (1.1) admits a non-trivial solution $f_{1}$ with $\lambda(f_{1})<\infty$, such that for every solution $f_{2}$ of (1.1) that is linearly independent with $f_{1}$ we have $\lambda(f_{2})=\infty$? The constructions in Theorem 1.6 and Theorem 1.7 together the comments above present such entire functions $A$ of finite order. With modifications of the methods there, we may construct more such coefficient functions $A$ with certain growth properties.
In the remainder of this paper, we provide the proof for Theorem 1.3, Theorem 1.4, Theorem 1.6 and Theorem 1.7 in Section 2-Section 5, respectively. In the following sections, we keep the permanent notation: $(f_{1},f_{2})$ is a normalized pair of solutions of (1.1), the quotient $F=f_{2}/f_{1}$ is a solution of $S(F)=2A$, and $E=f_{1}f_{2}=F/F^{\prime}$ is a solution of (1.8). The notations $\phi,\psi,\tau,F,G$ and so on in each section will represent different functions.
@@
For every pair $(m,2n)$ of two nonnegative integers $m$ and $n$, we consider the rational function $R_{m,n}(z)$ in (1.4). In this section, we fix two distinct pairs $(m_{1},2n_{1})$ and $(m_{2},2n_{2})$ and will sometimes omit them from notation. We shall assume that at least one of $m_{1}$ and $m_{2}$ is nonzero and at least one of $n_{1}$ and $n_{2}$ is nonzero. Note that, unlike in [4], $g_{m_{1},n_{1}}$ and $g_{m_{2},n_{2}}$ are locally univalent meromorphic functions. The process of gluing two different $g_{m,n}$s is almost the same as that in [4]: by restricting $g_{m_{1},n_{1}}$ to the lower half-plane $\check{H}$ and $g_{m_{2},n_{2}}$ to the upper half-plane $\hat{H}$, we then paste them together, using a quasiconformal surgery, producing a meromorphic function $F$. Then our Bank–Laine function will be $E=F/F^{\prime}$ and thus $A=B(E)/4$ as in (1.8). For, we will use some different notations from the ones used in [5].
@@
The asymptotic behaviors of the diffeomorphism $\phi$ satisfy:
@@
$\begin{split}c=\frac{1}{N}\log\left[\frac{\binom{m_{1}+2n_{1}}{m_{1}}}{\binom{m_{2}+2n_{2}}{m_{2}}}\frac{N!}{M!}\right]\quad\text{and}\quad\delta=\frac{1}{2}\min\{1,K\}.\end{split}$
@@
$\begin{split}\log g_{m_{1},n_{1}}(\phi(x))&=e^{\phi(x)}(1+O(\phi(x)e^{-\phi(x)}))=e^{\phi(x)}(1+O(\phi(x)e^{-2x/3})),\quad x\to\infty.\end{split}$
@@
(2.4) $\begin{split}\phi^{\prime}=\frac{g^{\prime}_{m_{2},n_{2}}}{g_{m_{2},n_{2}}}\frac{g_{m_{1},n_{1}}\circ\phi}{g^{\prime}_{m_{1},n_{1}}\circ\phi}=e^{Mx-N\phi(x)}\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}\frac{\sum_{i=0}^{m_{1}}A_{i}e^{i\phi(x)}}{\sum_{i=0}^{m_{2}}A_{i}e^{ix}}\frac{\sum_{j=0}^{2n_{1}}B_{j}e^{j\phi(x)}}{\sum_{j=0}^{2n_{2}}B_{j}e^{jx}},\end{split}$
@@
(2.5) $\begin{split}\phi^{\prime}&=e^{Mx-N\phi(x)+(m_{1}+2n_{1})\phi(x)-(m_{2}+2n_{2})x}\frac{1+O(e^{-\phi(x)})}{1+O(e^{-x})}\frac{1+O(e^{-\phi(x)})}{1+O(e^{-x})}\\ &=e^{O(e^{-x/2})}(1+O(e^{-\phi(x)})+O(e^{-x}))=1+O(e^{-x/2}),\quad x\to\infty.\end{split}$
@@
$\begin{split}\frac{\binom{m_{2}+2n_{2}}{m_{2}}}{\binom{m_{1}+2n_{1}}{m_{1}}}\frac{M!}{N!}e^{N\phi(x)-Mx}=1+O(e^{x})+O(e^{\phi(x)}),\quad x\to-\infty\end{split}$
@@
$\begin{split}\phi(x)=\frac{M}{N}x+\frac{1}{N}\log\left[\frac{\binom{m_{2}+2n_{2}}{m_{2}}}{\binom{m_{1}+2n_{1}}{m_{1}}}\frac{M!}{N!}\right]+O(e^{x})+O(e^{\phi(x)}),\quad x\to-\infty,\end{split}$
@@
$\begin{split}\phi^{\prime}(x)&=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{Mx-N\phi(x)}\frac{1+O(e^{\phi(x)})}{1+O(e^{x})}\frac{1+O(e^{\phi(x)})}{1+O(e^{x})}\\ &=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{Mx-N(Kx+c+O(e^{-\delta|x|/2}))}(1+O(e^{\phi(x)})+O(e^{x}))\\ &=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{-Nc+O(e^{-\delta|x|/2})}(1+O(e^{\phi(x)})+O(e^{x}))=K+O(e^{-\delta|x|/2}).\end{split}$
Now we begin to define a quasiregular function in the same way as in [5]. Though the functions $g_{m_{1},n_{1}}$ and $g_{m_{2},n_{2}}$ we use here may have some poles, the process of definition here is in many parts the same as that in [5]. For the convenience of calculating the number of zeros of the desired Bank–Laine functions in the final part of the proof and also of the applications of the arguments in the proof of Theorem 1.4, Theorem 1.6 and Theorem 1.7, below we shall keep as many as details.
@@
Note that $K\not=1$ under assumption. As shown in [5], $p$ maps $D$ onto the complement $G$ of a logarithmic spiral $\Gamma$, with
@@
The inverse map $h=p^{-1}$ is a conformal homeomorphism $h:G\to D$. Let $\Gamma^{\prime}=p(\mathbb{R}_{\geq 0})$. The two logarithmic spirals $\Gamma$ and $\Gamma^{\prime}$ divide the plane into two parts $\hat{G}$ and $\check{G}$, which are images under $p$ of the upper and lower half-planes, respectively.
@@
$V(z)=\left\{\begin{array}[]{ll}(g_{m_{2},n_{2}}\circ h)(z),&\quad z\in\hat{G},\\ (g_{m_{1},n_{1}}\circ h)(z),&\quad z\in\check{G}\end{array}\right.$
is meromorphic in $\hat{G}\cup\check{G}$ and has a jump discontinuity on $\Gamma$ and $\Gamma^{\prime}$. To remove this discontinuity, we consider the strip $\Pi=\{z:|\text{Im}\ z|<1\}$ and define a quasiconformal homeomorphism $\psi:\mathbb{C}\to\mathbb{C}$ to be the identity outside of $\Pi$ and for $y=\text{Im}\ z\in(-1,1)$,
(2.8) $\psi(x+iy)=\left\{\begin{array}[]{ll}\phi(x)+|y|(x-\phi(x))+iy,&\quad x\geq 0,\\ \phi(x/K)+|y|(x-\phi(x/K)+iy),&\quad x<0.\end{array}\right.$
Then $\psi$ commutes with the complex conjugation, and satisfies
(2.9) $\begin{split}\psi(x)=\phi(x),\quad x>0,\quad\text{and}\quad\psi(Kx)=\phi(x),\quad x<0.\end{split}$
The Jacobian matrix $D_{\psi}$ of $\psi$ is given for $x>0$ and $10<|y|<$ 1 by
$D_{\psi}(x+iy)=\left(\begin{array}[]{cc}\phi^{\prime}(x)+|y|(1-\phi(x))&\pm(x-\phi(x))\\ 0&1\\ \end{array}\right),$
@@
$D_{\psi}(x+iy)=\left(\begin{array}[]{cc}1&0\\ 0&1\\ \end{array}\right),\quad 0<|y|<1,\quad x\to\infty.$
@@
$D_{\psi}(x+iy)=\left(\begin{array}[]{cc}1&\mp c\\ 0&1\\ \end{array}\right),\quad 0<|y|<1,\quad x\to-\infty.$
We conclude that $\psi$ is quasiconformal in the plane.
@@
(2.10) $U(z)=\left\{\begin{array}[]{ll}(g_{m_{2},n_{2}}\circ h)(z),&\quad z\in\hat{G}\cup\Gamma\cup\Gamma^{\prime}\cup\{0\},\\ (g_{m_{1},n_{1}}\circ\psi\circ h)(z),&\quad z\in\check{G}.\end{array}\right.$
@@
It follows from (2.7) and (2.8) that $U$ is continuous on $\Gamma$ and $\Gamma^{\prime}$ and quasiregular in the plane. The existence theorem for solutions of the Beltrami equation [14, § V.1] shows that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ with the same Beltrami coefficient as $U$. The function $F=U\circ\tau^{-1}$ is then meromorphic. Moreover, as shown in [4], the Beltrami coefficient of $U$ (and hence of $\tau$) satisfies the hypotheses of the Techmüler–Wittich–Beliskii theorem [14, § V.6] (see also [17]). This theorem shows that $\tau$ is conformal at $\infty$ and may thus be normalized to satisfy
(2.11) $\begin{split}\tau(z)\sim z,\quad z\to\infty.\end{split}$
@@
We set $E=F/F^{\prime}$. As $F^{\prime}(z)\not=0$ for all $z\in\mathbb{C}$ by construction, $E$ is entire. As all poles and zeros of $F$ are simple, all residues of $F/F^{\prime}$ are equal to $1$ or $-1$, so $E^{\prime}(z)=1$ or $E^{\prime}(z)=-1$ at every zero of $E$, which implies the Bank–Laine property. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.
@@
We shall use a slightly different method from that in [4] to show that $E$, as well as the entire function $A$, has order $\rho=1/\text{Re}(\mu)$. Below we shall first estimate the numbers of poles and zeros of $F$ respectively. For two positive quantities $S_{1}(r)$ and $S_{2}(r)$, below we use the notation $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$.
@@
For the meromorphic function $g_{m_{2},n_{2}}$, if $m_{2}\not=0$, then the number of poles of $g_{m_{2},n_{2}}$ satisfies $n(r,\infty,g_{m_{2},n_{2}})\sim m_{2}r/\pi$ as $r\to\infty$. Thus, by the periodicity of $g_{m_{2},n_{2}}$, the number of poles of $g_{m_{2},n_{2}}$ in the upper half-pane $\hat{H}$ satisfies $\hat{n}(r,\infty,g_{m_{2},n_{2}})\sim m_{2}r/(2\pi)$ as $r\to\infty$. We note that the image of the circle $\{z:|z|=r\}$ under $h(z)=z^{1/\mu}$ is the part of the logarithmic spiral which connects two points on the negative real axis and intersects the positive real axis at $r^{1/\text{Re}(\mu)}$. By the definitions of $p$ and $h$ we may write $z=re^{i\theta}$ on the circle such that $z_{0}=re^{i\theta_{0}}$ is a point on $\Gamma^{\prime}$ and
$\begin{split}h(z)=\exp\left(\frac{1}{\mu}\log z\right)=\exp\left(\frac{(\text{Im}(\mu))\theta+(\text{Re}(\mu))\log r}{|\mu|^{2}}+i\frac{(\text{Re}(\mu))\theta-(\text{Im}(\mu))\log r}{|\mu|^{2}}\right).\end{split}$
@@
When $K>1$, noting that $\text{Re}(\mu)>0$ and $\text{Im}(\mu)<0$, we may write $-(\theta_{0}+\pi)\leq\theta\leq-(\theta_{0}-\pi)$, where $\theta_{0}$ satisfies $(\text{Re}(\mu))\theta_{0}=(\text{Im}(\mu))\log r$. A simple computation shows that
@@
Therefore, when $r$ is large, the number of poles of $g_{m_{2},n_{2}}\circ h$ in $\hat{G}$ satisfies
$\begin{split}c_{2}r^{1/\text{Re}(\mu)}(\log r)\leq\hat{n}(r,\infty,g_{m_{2},n_{2}}\circ h)\leq c_{1}r^{1/\text{Re}(\mu)}(\log r)\end{split}$
for two positive constants $c_{1}$ and $c_{2}$ both dependent on $K$. Similarly, if $m_{1}\not=0$, the number of poles of $g_{m_{1},n_{1}}\circ h$ in $\check{G}$ satisfies
(2.13) $\begin{split}c_{4}r^{1/\text{Re}(\mu)}(\log r)\leq\check{n}(r,\infty,g_{m_{1},n_{1}}\circ h)\leq c_{3}r^{1/\text{Re}(\mu)}(\log r)\end{split}$
for two positive constants $c_{3}$ and $c_{4}$ both dependent on $K$. By the definition of $\tau(x+iy)$ in (2.8) and the asymptotic behaviors for $\phi$ in (2.2) and (2.3), we see that the number of poles of $g_{m_{1},n_{1}}\circ\psi\circ h$ in $\check{G}$ satisfies
(2.14) $\begin{split}c_{6}r^{1/\text{Re}(\mu)}(\log r)\leq\check{n}(r,\infty,g_{m_{1},n_{1}}\circ\psi\circ h)\leq c_{5}r^{1/\text{Re}(\mu)}(\log r)\end{split}$
for two positive constants $c_{5}$ and $c_{6}$ both dependent on $K$. When $K<1$, we can also obtain two similar estimates as in (2.13) and (2.14). By summarizing the above estimates together with the relation in (2.11), we conclude that $n(r,\infty,F)\asymp r^{1/\text{Re}(\mu)}$. It follows that $N(r,\infty,F)\asymp r^{1/\text{Re}(\mu)}$. Similarly, when at least one of $n_{1}$ and $n_{2}$ is nonzero, for the zeros of $F$ we also obtain $n(r,0,F)\asymp r^{1/\text{Re}(\mu)}$ and thus $N(r,0,F)\asymp r^{1/\text{Re}(\mu)}$. Since $E=F/F^{\prime}$ and since the poles and zeros of $F$ are simple, we have
(2.15) $\begin{split}N\left(r,0,E\right)\asymp r^{1/\text{Re}(\mu)}(\log r).\end{split}$
@@
where $c$ is a suitably chosen positive constant and, together with the asymptotic behaviors for $\phi$ in (2.2) and (2.3) and also the definition $\psi(x+iy)$ in (2.9), that
$\begin{split}\log|g_{m_{1},n_{1}}\circ\psi\circ h|\leq c(L_{1}+2\varepsilon)|z|^{1/\text{Re}(\mu)}+\exp(c|z|^{1/\text{Re}(\mu)})\quad\text{for}\quad z\in G^{-}\quad\text{and}\quad|z|\not\in\Omega_{3}.\end{split}$
@@
as $|z|\to\infty$ outside a set $\Omega_{3}$ of finite linear measure. Together with the relation in (2.17), the lemma on the logarithmic derivative [10, Section 2.2] now implies that $E=F/F^{\prime}$ satisfies
@@
outside a set of finite linear measure. Thus, by (2.15) and the first main theorem of Nevanlinna theory [10, Section 1.3], the characteristic function $T(r,E)$ of $E$ satisfies
$\begin{split}T\left(r,E\right)\asymp r^{1/\text{Re}(\mu)}(\log r).\end{split}$